The denominator of a fraction exceeds the numeratorby 2. If 5 be added to the numerator the fraction increase by unity. The fraction is ?
Answer by Boreal(15235)

numerator is x
denominator is (x+2)
(x+5)/(x+2)=(x/(x+2))+1
But x/(x+2) +1=(x+x+2)/(x+2), a common denominator. That is (2x+2)/(x+2)
(x+5)/(x+2)=(2x+2)/(x+2)
The (x+2) cancel
x+5=2x+2
x=3
the original fraction is 3/5. ANSWER
add 5 to the numerator and it is 8/5, 1 3/5, and that is an increase by 1.